OpenStudy (anonymous):

Solve each formula for the indicated letter. Assume that all variables represent nonnegative numbers. W= √1/LC, for L (An electricity formula)

OpenStudy (saifoo.khan):

\[W = \sqrt{\frac{1}{LC}}\] take square on both sides, \[W^2 = \frac{1}{LC}\] \[L = \frac{1}{LW^2}\]

OpenStudy (saifoo.khan):

Oh man! Typo in the last one, it's \[L = \frac{1}{W^2}\]

OpenStudy (anonymous):

is that the whole thing?

OpenStudy (saifoo.khan):

i showed you steps.